A barometer is filled with a cylindrical column of mercury that is 76.0 cm high and 1.000 cm in diameter. If the density of mercury is 13.53 g/cm3, what is the mass of mercury in the column? (NOTE: Vol. of cylinder = Π x radius2 x height )
1
Expert's answer
2017-09-14T02:43:06-0400
V = π∙(d/2)^2 ∙h = π ∙ 0.5^2 ∙ 76 = 59.69 (cm^3) - the volume of cylinder
m = V∙ρ = 59.69 ∙ 13.53 = 807.6 (g) - the mass of mercury
Answer: the mass of mercury is 807.6 g